OpenStudy (anonymous):

Evaluate the limit, if it exist lim as x approaches 1, (x^3-1)/(x^2-1).

OpenStudy (anonymous):

hint: exponents!

OpenStudy (anonymous):

please show step by step thanks

OpenStudy (amistre64):

3x/2x = 3/2

OpenStudy (anonymous):

i dont understand

OpenStudy (amistre64):

it the general hospital rule...

OpenStudy (anonymous):

ok factor top and bottom \[\frac{x^3-1}{x^2-1}=\frac{(x-1)(x^2+x+1)}{(x+1)(x-1)}\]

OpenStudy (anonymous):

now cancel the x - 1 to get \[\frac{x^2+x+1}{x+1}\]

OpenStudy (anonymous):

replace x by 1, get \[\frac{3}{2}\]

OpenStudy (anonymous):

oooooh ok i tried the other way and thought that you could factor it out...im still trying to understand where the x^2+x +1 came from

OpenStudy (anonymous):

this kind of problem tends to come before differentiation and therefore before l'hopital's rule

OpenStudy (anonymous):

difference of two cubes! \[a^3-b^3=(a-b)(a^2+ab+b^2)\]

OpenStudy (anonymous):

in your case a = x and b = 1 to give \[x^3-1=(x-1)(x^2+x+1)\]

OpenStudy (anonymous):

so basically you were factoring it out

OpenStudy (anonymous):

exactly

OpenStudy (anonymous):

ok cool i knew that what the steps were...well thanks again for the help

OpenStudy (anonymous):

btw if you replace x by 1 in the numerator you get 0 yes? this means the numerator MUST factor as (x-1) times something.

OpenStudy (anonymous):

yea.. thats what i did but somehow did it wrong

OpenStudy (anonymous):

it is not a miracle you can factor, it is a consequence of the fact that you get 0 if you replace x by 1. same as in the denominator

OpenStudy (anonymous):

yea thats exactly what i did so i figure theres something more to it

OpenStudy (anonymous):

thanks again your a life saver